OpenStudy (anonymous):

(x+2)^2 +11(x+2)+24 =0

OpenStudy (anonymous):

x^2+15x+50 x=-10, x=-5

OpenStudy (anonymous):

(x+2)^2+11(x+2)+24 x^2+4x+4+11x+22+24 x^2+15x+50 (x+5)(x+10) x=-5,-10

myininaya (myininaya):

rickie read your previuous thread i answer this question

OpenStudy (anonymous):

k thanks so much